# Clairaut anti-invariant Riemannian map from/to Kähler manifolds admitting Ricci solitons

### v1

@@ Abstract
The aim of this article is to study the Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons. We find the curvature relations and calculate the Ricci tensor under different conditions. We obtain conditions for the range and kernel of these maps to be Einstein. Next, we find the scalar curvature for range. Finally, we give some non-trivial examples of Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons.
@@
Keywords: Kähler manifolds, Riemannian maps, Clairaut Riemannian maps, anti-invariant Riemannian maps, Ricci solitons.
@@ 1 Introduction
The submersions and immersions play an important role in the study of smooth maps between smooth manifolds. In the theory of Riemannian manifolds, Riemannian maps were introduced by Fisher [4] in 1992 as a generalization of isometric immersion and Riemannian submersion. The notable property of Riemannian maps is that these maps satisfy the generalized eikonal equation, which connects the geometric and physical optics.
The geometry of Riemannian submersions was discussed by Falcitelli et al. [3]. As a generalization of holomorphic Riemannian submersions, Şahin [15] introduced holomorphic Riemannian maps. Further, as a generalization of anti-invariant Riemannian submersions, Şahin [14] introduced anti-invariant Riemannian maps from almost Hermitian manifolds to Riemannian manifolds.
The idea of Clairaut theorem is based on geodesics on the surface of revolution. This theorem states that for any geodesic $\gamma$($\gamma:I\subset R\rightarrow S$) on a surface of revolution, $rsin\theta$ is constant along $\gamma$, where $\theta(s)$ is the angle between $\gamma(s)$ and the meridian curve through $\gamma(s)$, $s\in I$ and $r:S\rightarrow R^{+}$ is the distance of a point of $S$ from the axis of rotation. Bishop [1] generalized this idea on submersion theory and defined Clairaut submersion. In [18], Şahin gave the idea of Clairaut Riemannian maps, in which he obtained the necessary and sufficient conditions for a Riemannian map to be a Clairaut Riemannian map. Recently, Meena et al. [7, 22] have studied Clairaut anti-invariant Riemannian maps from/to Kähler manifolds.
@@
A smooth vector field $Z$ on a Riemannian manifold is called a conformal vector field if the Lie derivative of the metric g with respect to X can be expressed as a multiple of a function to the metric g, and this function is called the conformal factor of the conformal vector field $Z$ [2].
@@
$\frac{1}{2}(L_{\xi}g)(X,Y)+\alpha Ric(X,Y)+\lambda g_{M}(X,Y)=0,$ (1.1)
where $\alpha,\lambda\in\operatorname{\mathbb{R}},$ $\xi$ is a potential vector field, $L_{\xi}g$ denotes the Lie derivative of metric tensor $g$ with respect to $\xi$, Ric denotes the Ricci curvature.
The $\alpha$ Ricci soliton is called shrinking, steady or expanding according as $\lambda<0,\lambda=0$ or $\lambda>0,$ respectively. If $\alpha=1,$ then (1.1) reduces to Ricci soliton. If $\xi$ is gradient of some function $f$ which is defined on $M$, i.e., $\xi=gradf$, $f\in C^{\infty}(M)$, then from (1.1), we get
$Hess^{f}(X,Y)+Ric(X,Y)+\lambda g(X,Y)=0,$
where $Hess^{f}$ is hessian form. Then $(M,g_{M},\nabla f,\lambda)$ is called gradient Ricci soliton and the function f is called potential function.
Pigola et al. [11] gave the idea of almost Ricci soliton by using $\lambda$ as a variable function rather than constant function. In this case $(M,g_{M},\xi,\lambda)$ is called an almost Ricci soliton.
Meriç and Kiliç [8] study about Riemannian submersions whose total manifolds admit a Ricci soliton. Yadav and meena [21] have studied Riemannian maps whose total admit a Ricci soliton. In [7, 23], Meena and yadav have studied Clairaut Riemannian maps whose total and base manifolds admit a Ricci soliton. Watson [20] studies Hermitian submersion whose total manifold admits Ricci soliton.
In this article, we study Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons. The article is divided into four sections. In section 2, we recall some definitions and results which are useful to current study. In section 3, we study Clairaut anti-invariant Rieannian maps from a Kähler manifold admitting Ricci soliton to a Riemannian manifold and give non trivial examples for the existence. In section 4, we study Clairaut anti-invariant Riemannian maps from a Riemannian manifold to a Kähler manifold admitting Ricci soliton and construct some non-trivial examples.
@@ 2 Preliminaries
Let $F$ be a smooth map between two Riemannian manifolds $(M,g_{M})$ and $(N,g_{N})$ of dimension $m,n,$ respectively. Let $\mathcal{V}_{p}=kerF_{*p}$ at $p\in M,$ denotes vertical distribution or kernel space of $F_{*}$ and $\mathcal{H}_{p}=(kerF_{*p})^{\perp}$ in $T_{p}M$ is the orthogonal complementary space of $\mathcal{V}_{p}$. Then the tangent space $T_{p}M$ at $p\in M$ has the decomposition,
@@
Let the range of $F_{*}$ be denoted by $rangeF_{*p}$ at $p\in M$, and $(rangeF_{*p})^{\perp}$ be the orthogonal complementary space of $rangeF_{*p}$ in the tangent space $T_{F(p)}N$ of $N$ at $F(p)\in N$. Since $rankF\leq\min\{m,n\}$, therefore we get, $(rangeF_{*p})^{\perp}\neq\{0\}$. Thus, the tangent space $T_{F(p)}N$ of $N$ at $F(p)\in N$ has the following decomposition:
@@
A smooth map $F$ is said to be a Riemannian map at $p\in M$ if the horizontal restriction $F^{h}_{p}:(kerF_{*p})^{\perp}\rightarrow(rangeF_{*p})$ is a linear isometry between the inner product spaces $((kerF_{*p})^{\perp},g_{N}{(p)}|(kerF_{*p})^{\perp})$ and $((rangeF_{*p},g_{N}(q)|(rangeF_{*p})),q=F(p)$. In other words, $F_{*}$ satisfies the equation [17]
@@
For Riemannian submersion, O’Neill [10] defined fundamental tensor fields $\mathcal{T}$ and $\mathcal{A}$ by
$A_{E}{F}=\mathcal{H}\nabla_{HE}VF+\mathcal{V}\nabla_{HE}HF,\$ (2.2)
$T_{E}{F}=\mathcal{H}\nabla_{VE}VF+\mathcal{V}\nabla_{VE}HF,\$ (2.3)
where $E,F\in\Gamma(M),\nabla$ is the Levi-Civita connection of $g_{M}$ and $\mathcal{V},\mathcal{H}$ denote the projections to vertical subbundle and horizontal subbundle, respectively. For any $E\in\Gamma(TM)$, $T_{E}$ and $A_{E}$ are skewsymmetric operators on $(\Gamma(TM),g)$ reversing the horizontal and the vertical distributions. One can check easily that $\mathcal{T}$ is vertical, i.e., $\mathcal{T}_{E}=T_{\mathcal{V}E}$, and $\mathcal{A}$ is horizontal, i.e., $\mathcal{A}=A_{HE}$.
We observe that the tensor field $T$ is symmetric for vertical vector fields, i.e.,
$T_{U}W=T_{W}U~{}~{}\forall~{}~{}U,W\in\Gamma(kerF_{*}),\$
and tensor field $A$ is anti-symetric for horizontal vector fields i.e.,
$A_{X}Y=-A_{Y}X=\frac{1}{2}\mathcal{V}[X,Y]~{}~{}\forall~{}~{}X,Y\in\Gamma(kerF_{*})^{\perp}.\$
Using equation (2.2) and (2.3), we have the following lemma.
@@
$\nabla_{X}V=\mathcal{A}_{X}V+\mathcal{V}\nabla_{X}V,$
$\nabla_{X}Y=\mathcal{H}\nabla_{X}Y+\mathcal{A}_{X}Y.$
@@
For any vector field $X$ on $M$ and any section $V$ of $(rangeF_{*})^{\perp}$, we define the connection $\nabla^{F\perp}$ as the orthogonal projection of $\nabla_{X}V$ on $(rangeF_{*})^{\perp}.$ One can easily see that $\nabla^{F\perp}$ is a linear connection on $(rangeF_{*})^{\perp}$ with $\nabla^{F\perp}g_{N}=0.$ For a Riemannian map $F$, we have [16]
$\nabla^{N}_{F_{*}X}V=-S_{V}{F_{*}X}+\nabla^{F\perp}_{X}V,$ (2.4)
where $S_{V}{F_{*}X}$ is the tangential component and $\nabla^{F\perp}_{X}V$ is the orthogonal component of $\nabla^{N}_{F_{*}X}V$. It can be easily seen that $\nabla^{N}_{F_{*}X}V$ is obtained from the pullback connection of $\nabla^{N}$. Thus, at $p\in M$, we have $\nabla^{N}_{F_{*}X}V(p)\in T_{F(p)}N,S_{V}{F_{*}X}\in F_{*p}(T_{p}M)$ and $\nabla^{F\perp}_{X}V\in(F_{*p}(T_{p}M))^{\perp}$. It follows that $S_{V}{F_{*}X}$ is bilinear in $V$, $F_{*}X$ and $S_{V}{F_{*}X}$ at $p$ depends only on $V_{p}$ and $F_{*p}X_{p}$.
@@
$g_{N}(S_{V}{F_{*}X,{F_{*}Y}})=g_{N}\big{(}V,(\nabla F_{*})(X,Y)\big{)}\$
for all X, Y $\in\Gamma(kerF_{*})^{\perp}~{}and~{}V\in\Gamma(rangeF_{*})^{\perp}.$ Since $\nabla F_{*}$ is symmetric, it follows that $S_{V}$ is a symmetric linear transformation of $rangeF_{*}$.
@@
$(\nabla F_{*})(X,Y)=\nabla^{N_{F}}_{X}{F_{*}}Y-F_{*}(\nabla^{M}_{X}Y),$
@@
Theorem 2.1.
[7] Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a Riemannian map between Riemannian manifolds such that $(rangeF_{*})^{\perp}$ is totally geodesic and $rangeF_{*}$ is connected, let $\beta$ and $\eta=Fo\beta$ be geodesics on $M$ and $N,$ respectively.Then $F$ is a Clairaut Riemannian map with $\tilde{s}=e^{g}$ if and only if any one of the following conditions holds:
  • (i)
    $S_{V}F_{*}X=-V(g)F_{*}X$, where $F_{*}X\in\Gamma(rangeF_{*})$ and $V\in\Gamma(rangeF_{*})^{\perp}$ are components of $\dot{\eta}(t)$.
  • (ii)
    $F$ is an umbilical map, and has $H=-\nabla^{N}g$ , where $g$ is a smooth function on $N$ and $H$ is the mean curvature vector field of $rangeF_{*}$.
Definition 2.1.
[7] A Riemannian map $F:(M,g_{M})\rightarrow(N,g_{N})$ between Riemannian manifolds is called a Clairaut Riemannian map if there is a function $s:N\rightarrow R^{+}$ such that for every geodesic $\eta$ on $N,$ the function $(so\eta)sin\omega(t)$ is constant, where $F_{*}X\in\Gamma(rangeF_{*})$ and $U\in\Gamma(rangeF_{*})^{\perp}$ are the vertical and horizontal components of $\dot{\eta}(t)$, and $\omega(t)$ is the angle between $\dot{\eta}(t)$ and $U$ for all t.
@@
$(\nabla F_{*})(X,Y)=g_{M}(X,Y)H$
for all $X,Y\in\Gamma(kerF_{*})^{\perp}$ and $H$ is nowhere zero mean curvature vector field on $(rangeF_{*})^{\perp}$.
@@
$g_{M}(JX,JY)=g_{M}(X,Y)$ (2.5)
for all $X,Y\in\Gamma(TM).$ An almost Hermitian manifold $M$ is called a Kähler manifold if
$(\nabla_{X}J)Y=0\$ (2.6)
for all $X,Y\in\Gamma(TM)$, where $\nabla$ is Levi-Civita connection on $M$.
@@
[25] For a Kähler manifold $M$, follwing properties are satisfied:
  • (i)
    R(X, Y) JZ = JR(X, Y)Z,
  • (ii)
    R(JX, JY) = R(X, Y),
  • (iii)
    Ric(X, Y) = $\frac{1}{2}\{Tr.JoR(X,JY)\},QJ=JQ,$
  • (iv)
    Ric(JX, JY) = Ric(X, Y),
where $X,Y,Z\in\Gamma(TM)$, $Q$ is the Ricci operator and $R$ is Riemannian curvature tensor.
@@
Next, let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be an anti-invariant Riemannian map from a Kähler manifold to a Riemannian manifold. Then, for any $X\in\Gamma((kerF_{*})^{\perp}),$ we have
@@
where $BX\in\Gamma(kerF_{*})$ and $CX\in\Gamma(\mu).$

### v2

# Clairaut anti-invariant Riemannian maps with Kähler and Ricci soliton structures

@@ Abstract
The aim of this article is to explore the Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons. We find the curvature relations and calculate the Ricci tensor under different conditions. We discuss the condition under which range space becomes $\alpha$-Ricci soliton. We obtain conditions for the range and kernel spaces of these maps to be Einstein. Next, we find the scalar curvature for range space. Further, we give the relation between Ricci curvature and Lie derivative under these maps. Moreover, we find the condition for a vertical potential vector field on target manifold to be conformal vector field on range space of these maps. Finally, we give non-trivial examples of such maps.
@@
Keywords: Kähler manifolds, Riemannian maps, Clairaut Riemannian maps, anti-invariant Riemannian maps, Ricci solitons
@@ 1 Introduction
The submersions and immersions play an important role in the study of smooth maps between smooth manifolds. In the theory of Riemannian manifolds, Riemannian maps were introduced by Fischer [4] in 1992 as a generalization of isometric immersion and Riemannian submersion. The notable property of Riemannian maps is that these maps satisfy the generalized eikonal equation, which connects the geometric and physical optics.
The geometry of Riemannian submersions was discussed by Falcitelli et al. [3]. As a generalization of holomorphic Riemannian submersions, Şahin [23] introduced holomorphic Riemannian maps. As a generalization of anti-invariant Riemannian submersions, Şahin [22] introduced anti-invariant Riemannian maps from almost Hermitian manifolds to Riemannian manifolds.
The idea of Clairaut’s Theorem is based on geodesics on the surface of revolution. This theorem states that for any geodesic $\gamma:I\subset\mathbb{R}\rightarrow S$ on a surface of revolution, $r\sin\theta$ is constant along $\gamma$, where $\theta(s)$ is the angle between $\gamma(s)$ and the meridian curve through $\gamma(s)$, $s\in I$ and $r:S\rightarrow\mathbb{R}^{+}$ is the distance of a point of $S$ from the axis of rotation. The authors of [1, 10] generalized this idea on submersion theory and defined Clairaut submersions. Further, this idea has been generalised to Riemannian maps as necessary and sufficient conditions for Riemannian maps to be a Clairaut Riemannian maps obtained in [26, 11, 13]. Recently, Clairaut anti-invariant Riemannian maps from/to Kähler manifolds have been studied in [11, 30].
@@
A smooth vector field $X^{\prime}$ on a Riemannian manifold $(M,g_{M})$ is called a conformal vector field if the Lie derivative of the metric $g_{M}$ with respect to $X^{\prime}$ can be expressed as a multiple of a function to the metric $g_{M}$, and this function is called the conformal factor of the conformal vector field $X^{\prime}$ [2].
@@
$\frac{1}{2}(L_{\xi}g_{M})(X,Y)+\alpha Ric(X,Y)+\lambda g_{M}(X,Y)=0,$ (1.1)
where $\alpha,\lambda\in\operatorname{\mathbb{R}},$ $\xi$ is a potential vector field, $L_{\xi}g_{M}$ denotes the Lie derivative of metric tensor $g_{M}$ with respect to $\xi$, Ric denotes the Ricci curvature. The $\alpha$ Ricci soliton is called shrinking, steady or expanding according to $\lambda<0,\lambda=0$ or $\lambda>0,$ respectively. If $\alpha=1,$ then (1.1) reduces to Ricci soliton. If $\xi$ is gradient of some function $f$ which is defined on $M$, i.e., $\xi=gradf$, $f\in C^{\infty}(M)$, then from (1.1), we get
$Hess^{f}(X,Y)+Ric(X,Y)+\lambda g_{M}(X,Y)=0,$
where $Hess^{f}$ is hessian form. Then $(M,g_{M},\nabla f,\lambda)$ is called gradient Ricci soliton and the function f is called potential function. Pigola et al. [18] gave the idea of almost Ricci soliton by using $\lambda$ as a variable function rather than constant function. In this case $(M,g_{M},\xi,\lambda)$ is called an almost Ricci soliton.
The authors of [14, 15, 12] study about submersions whose total manifolds admit a Ricci soliton. Yadav, Meena [29, 32] and Gupta et al. [7] have studied Riemannian maps whose total (base) manifold admit a Ricci soliton. In [11, 31, 13], Meena and Yadav have studied Clairaut Riemannian maps whose total (base) manifolds admit a Ricci soliton. Watson [28] introduced an almost Hermitian submersion and Gündüzalp [6] studied an almost Hermitian submersion [34] whose total manifold admits Ricci soliton.
In this paper, we explore Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons. The paper is divided into five sections. In section 2, we recall some definitions and results which are useful to current study. In section 3, we study Clairaut anti-invariant Riemannian maps from a Kähler manifold admitting Ricci soliton to a Riemannian manifold and give non trivial examples for the existence. In section 4, we study Clairaut anti-invariant Riemannian maps from a Riemannian manifold to a Kähler manifold admitting Ricci soliton and construct some non-trivial examples. In section 5, we give scope for further studies based on current study.
@@ 2 Preliminaries
Let $F$ be a smooth map between two Riemannian manifolds $(M,g_{M})$ and $(N,g_{N})$ of dimension $m,n,$ respectively. Let $\mathcal{V}_{p}=kerF_{*p}$ at $p\in M,$ denotes vertical distribution or kernel space of $F_{*}$ and $\mathcal{H}_{p}=(kerF_{*p})^{\perp}$ in $T_{p}M$ is the orthogonal complementary space of $\mathcal{V}_{p}$. Then, we decompose the tangent space $T_{p}M$ at $p\in M$ in the following manner.
@@
Let the range of $F_{*}$ be denoted by $rangeF_{*p}$ at $p\in M$, and $(rangeF_{*p})^{\perp}$ be the orthogonal complementary space of $rangeF_{*p}$ in the tangent space $T_{F(p)}N$ of $N$ at $F(p)\in N$. Since $rankF<\min\{m,n\}$, we get $(rangeF_{*p})^{\perp}\neq\{0\}$. Thus, the tangent space $T_{F(p)}N$ of $N$ at $F(p)\in N$ has the following decomposition:
@@
A smooth map $F$ is said to be a Riemannian map at $p\in M$ if the horizontal restriction $F^{h}_{p}:(kerF_{*p})^{\perp}\rightarrow(rangeF_{*p})$ is a linear isometry between the inner product spaces $((kerF_{*p})^{\perp},g_{M}{(p)}|(kerF_{*p})^{\perp})$ and $((rangeF_{*p},g_{N}(q)|(rangeF_{*p})),q=F(p)$. In other words, $F_{*}$ satisfies the equation [25]
@@
For Riemannian submersion, O’Neill [17] defined fundamental tensor fields $T$ and $A$ by
$A_{X^{\prime}}{Y^{\prime}}=\mathcal{H}\nabla^{M}_{\mathcal{H}X^{\prime}}\mathcal{V}Y^{\prime}+\mathcal{V}\nabla^{M}_{\mathcal{H}X^{\prime}}\mathcal{H}Y^{\prime},\$ (2.2)
$T_{X^{\prime}}{Y^{\prime}}=\mathcal{H}\nabla^{M}_{\mathcal{V}X^{\prime}}\mathcal{V}Y^{\prime}+\mathcal{V}\nabla^{M}_{\mathcal{V}X^{\prime}}\mathcal{H}Y^{\prime},\$ (2.3)
where $X^{\prime},Y^{\prime}\in\Gamma(TM),\nabla^{M}$ is the Levi-Civita connection of $g_{M}$ and $\mathcal{V},\mathcal{H}$ denote the projections to vertical subbundle and horizontal subbundle, respectively. For any $X^{\prime}\in\Gamma(TM)$, $T_{X^{\prime}}$ and $A_{X^{\prime}}$ are skewsymmetric operators on $(\Gamma(TM),g_{M})$ reversing the horizontal and the vertical distributions. One can check easily that $T$ is vertical, i.e., $T_{X^{\prime}}=T_{\mathcal{V}X^{\prime}}$, and $A$ is horizontal, i.e., $A=A_{\mathcal{H}X^{\prime}}$.
We observe that the tensor field $T$ is symmetric for vertical vector fields, and tensor field $A$ is anti-symmetric for horizontal vector fields. Using equation (2.2) and (2.3), we have the following Lemma.
@@
$\nabla_{X}V=A_{X}V+\mathcal{V}\nabla_{X}V,$
$\nabla_{X}Y=\mathcal{H}\nabla_{X}Y+A_{X}Y.$
@@
For any vector field $X^{\prime}$ on $M$ and any section $D$ of $(rangeF_{*})^{\perp}$, we define the connection $\nabla^{F\perp}$ as the orthogonal projection of $\nabla_{X^{\prime}}D$ on $(rangeF_{*})^{\perp}.$ One can easily see that $\nabla^{F\perp}$ is a linear connection on $(rangeF_{*})^{\perp}$ with $\nabla^{F\perp}g_{N}=0.$ For a Riemannian map $F$, we have [24]
$\nabla^{N}_{F_{*}X}D=-S_{D}{F_{*}X}+\nabla^{F\perp}_{X}D,$ (2.4)
where $S_{D}{F_{*}X}$ is the tangential component and $\nabla^{F\perp}_{X}D$ is the orthogonal component of $\nabla^{N}_{F_{*}X}D$. It can be easily seen that $\nabla^{N}_{F_{*}X}D$ is obtained from the pullback connection of $\nabla^{N}$. Thus, at $p\in M$, we have $\nabla^{N}_{F_{*}X}D(p)\in T_{F(p)}N,S_{D}{F_{*}X}\in F_{*p}(T_{p}M)$ and $\nabla^{F\perp}_{X}D\in(F_{*p}(T_{p}M))^{\perp}$. It follows that $S_{D}{F_{*}X}$ is bilinear in $D$, $F_{*}X$ and $S_{D}{F_{*}X}$ at $p$ depends only on $D_{p}$ and $F_{*p}X_{p}$.
@@
$g_{N}(S_{D}{F_{*}X,{F_{*}Y}})=g_{N}\big{(}D,(\nabla F_{*})(X,Y)\big{)}\$
for all X, Y $\in\Gamma(kerF_{*})^{\perp}~{}and~{}V\in\Gamma(rangeF_{*})^{\perp}.$ Since $\nabla F_{*}$ is symmetric, it follows that $S_{D}$ is a symmetric linear transformation of $rangeF_{*}$.
@@
$(\nabla F_{*})(X^{\prime},Y^{\prime})=\nabla^{N_{F}}_{X^{\prime}}{F_{*}}Y^{\prime}-F_{*}(\nabla^{M}_{X^{\prime}}Y^{\prime}),$
@@
$(\nabla F_{*})(X,Y)=g_{M}(X,Y)H^{\prime}$
for all $X,Y\in\Gamma(kerF_{*})^{\perp}$ and $H^{\prime}$ is nowhere zero mean curvature vector field on $(rangeF_{*})^{\perp}$.
@@
$g_{M}(JX^{\prime},JY^{\prime})=g_{M}(X^{\prime},Y^{\prime})$ (2.5)
for all $X^{\prime},Y^{\prime}\in\Gamma(TM).$ An almost Hermitian manifold $M$ is called a Kähler manifold if
$(\nabla^{M}_{X^{\prime}}J)Y^{\prime}=0\$ (2.6)
for all $X^{\prime},Y^{\prime}\in\Gamma(TM)$, where $\nabla^{M}$ is Levi-Civita connection on $M$.
2.1 Some results on Clairaut anti-invariant Riemannian maps from Kähler manifolds
In this subsection, we recall some definitions and results which will be used in section 3. The first definition deals with Clairaut Riemannian map for source manifold which was defined by Şahin.
Definition 2.1.
[26] A Riemannian map $F:(M,g_{M})\rightarrow(N,g_{N})$ between Riemannian manifolds is called a Clairaut Riemannian map if there is a function $\tilde{r}:M\rightarrow\mathbb{R}^{+}$ such that for every geodesic, making angles $\theta$ with the horizontal subspaces, $\tilde{r}\sin\theta$ is constant.
@@ 2.1 Some results on Clairaut anti-invariant Riemannian maps from Kähler manifolds
Next, let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be an anti-invariant Riemannian map from a Kähler manifold to a Riemannian manifold. Then, for any $X\in\Gamma(kerF_{*})^{\perp},$ we have
@@
where $BX\in\Gamma(kerF_{*})$ and $CX\in\Gamma(\mu).$ Note that if $\mu=0,$ then $F$ is called a Lagrangian Riemannian map.
Now, we recall the definition of Clairaut anti-invariant Riemannian map for source manifold.
Definition 2.3.
[30] An anti-invariant Riemannian map from Kähler manifold to a Riemannian manifold is called Clairaut anti-invariant Riemannian map if it satisfies the definition of Clairaut Riemannian map for source manifold.
2.2 Some results on Clairaut anti-invariant Riemannian maps to Kähler manifolds
In this subsection, we recall some definitions and results which will be used in section 4. The first definition is about Clairaut Riemannian map for target manifold.
Definition 2.4.
[11] A Riemannian map $F:(M,g_{M})\rightarrow(N,g_{N})$ between Riemannian manifolds is called a Clairaut Riemannian map if there is a function $\tilde{s}:N\rightarrow\mathbb{R}^{+}$ such that for every geodesic $\sigma$ on $N,$ the function $(\tilde{s}o\sigma)sin\omega(t)$ is constant, where $F_{*}X\in\Gamma(rangeF_{*})$ and $D\in\Gamma(rangeF_{*})^{\perp}$ are the vertical and horizontal components of $\dot{\sigma}(t)$, and $\omega(t)$ is the angle between $\dot{\sigma}(t)$ and $D$ for all t.